OpenStudy (anonymous):

Solve for m. -3 + m/9=10 (/)means divided by.

sam (.sam.):

\[-3+\frac{m}{9}=10\] Multiply all terms by 9 \[-27+m=90\] m=117

OpenStudy (anonymous):

Multiply both sides by 9 to get rid of the fraction. That leaves you with \(\ \huge -27+m=90 \). Now, add 27 to isolate the variable. That leaves you with \(\ \huge m=90 \).

OpenStudy (anonymous):

Wait, sorry \(\ \huge 117 \). Forgot to add as I was typing!

sam (.sam.):

\[\frac{-3+m}{9}=10\] Multiply both sides by 9 \[-3+m=90\] m=93

OpenStudy (anonymous):